17. [5 points] Given the following statements, mark those correct statements as True and mark those incorrect statements as False (7n2 3 logn +22) True False O(ns 3m2 7logn 20 True False 5n 3 logn 1 O(nlogn) True False 7nlogn 3n. O(11n logn. 7) True False 2n2 3nlogne 7n 4logna 1 O(2n 3n 7) True False
Answer #1

n2 = O(7n2 + 3logn + 22) is correct since nighest power of n is same in LHS and RHS

2n = O(n3 +3n2 + 7logn + 2) is not true since highesest power of LHS is 2n bt of RHS is n2

5n + 3logn + 1 = O(nlogn) is also not true ssince highest power of n is 1 in LHS, and in RHS its nlogn and are not equal

7nlogn + 3n = O(11n + 5logn + 7) you can see that on ih RHS it is n + logn which is not equal to nlogn in left

2n2+3nlogn+7n+4logn + 1 = O(2n + 3n + 7) is again false since LHS has highest power n2 and of RHS it is 2n